Net Present Value Method, Present Value Index, and Analysis for a service company

Continental Railroad Company is evaluating three capital investment proposals by using the net present value method. Relevant data related to the proposals are summarized as follows:

Maintenance
Equipment
Ramp
Facilities
Computer
Network
Amount to be invested $568,602 $336,893 $171,869
Annual net cash flows:
Year 1 277,000 186,000 119,000
Year 2 258,000 167,000 82,000
Year 3 235,000 149,000 60,000
Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162

Required:

1. Assuming that the desired rate of return is 20%, prepare a net present value analysis for each proposal. Use the present value of $1 table above. If required, use the minus sign to indicate a negative net present value. If required, round to the nearest dollar.

Maintenance Equipment Ramp Facilities Computer Network
Total present value of net cash flow $ $ $
Amount to be invested
Net present value $ $ $

2. Determine a present value index for each proposal. If required, round your answers to two decimal places.

Present Value Index
Maintenance Equipment
Ramp Facilities
Computer Network
Answer #1
1
Maintenance Equipment Ramp Facilities Computer Network
Total present value of net cash flow 545858 357107 190775
Amount to be invested 568602 336893 171869
Net present value -22744 20214 18906
2
Present Value Index
Maintenance Equipment 0.96 =545858/568602
Ramp Facilities 1.06 =357107/336893
Computer Network 1.11 =190775/171869
Workings:
Maintenance Equipment:
Year PV factor Cash flows Present value
1 0.833 277000 230741
2 0.694 258000 179052
3 0.579 235000 136065
Total 545858
Ramp Facilities :
Year PV factor Cash flows Present value
1 0.833 186000 154938
2 0.694 167000 115898
3 0.579 149000 86271
Total 357107
Computer Network:
Year PV factor Cash flows Present value
1 0.833 119000 99127
2 0.694 82000 56908
3 0.579 60000 34740
Total 190775